OpenStudy (anonymous):

a falling stone takes 0.30s to travel past a window 2.2m tall. (a) from what height above the top of the window did the stone fall? (b) this was a second floor window, and the stone was dropped from the roof. How many floors tall is the building? (c) what are the signs of the stone's velocity and acceleration as the stone passes the window?

OpenStudy (amistre64):

calculating for an initial velocity might prove beneficial h = -1/2 g t^2 + vo t it travels 2.2m in .30 sec sooo 3.3 = -1/2 g (.3)^2 + .3 vo solve for vo so that we know the velocity that it starts to hit the window at knowing the velocity it is traveling allows us to find the initial height since we know how long it has to fall to reach that velocity vo = gt therefore time, t, falling is vo/g solving height again as: h = -1/2 g t^2, such that t=vo/g h = -vo^2/2g

OpenStudy (amistre64):

i cant ever recall the v = sqrt(2gh) formula so i always have to work into it i spose